Costs associated with the manufacture of miniature high-sensitivity piezoresistive pressure transducers is $74,000 per year. A clever industrial engineer found that by spending $16,000 now to reconfigure the production line and reprogram two of the robotic arms, the cost will go down to $55,000 next year and $57,000 in years 2 through 5. Using an interest rate of 8% per year, determine the present worth of the savings due to the reconfiguration.
Answer #1

savings in yr 1 = 74000 - 55000 = 19000

Savings in yr 2 to 5 = 74000 - 57000 = 17000

i = 8%

NPW of savings = -16000 + 17000*(P/A,8%,5) + 2000*(P/F,8%,1)

= -16000 + 17000*3.99271 + 2000*0.925926

= 53727.92